OpenStudy (anonymous):

simplify:-sec^2x/cot^2x+1

hero (hero):

\[\space \space \space -\sec^2x \div (\cot^2x + 1)\] \[=-\frac{1}{\cos^2x} \div \left(\frac{\cos^2x}{\sin^2x} + 1\right)\] \[=-\frac{1}{\cos^2x} \div \left(\frac{\cos^2x}{\sin^2x} + \frac{\sin^2x}{\sin^2x}\right)\] \[=-\frac{1}{\cos^2x} \div \left(\frac{1}{\sin^2x} \right)\] \[=-\frac{1}{\cos^2x} \times \left(\frac{\sin^2x}{1} \right)\] \[=-\frac{\sin^2x}{\cos^2x}\] \[=-\tan^2x\]

OpenStudy (anonymous):

in my txbook the answer is tan^2x

hero (hero):

There's a negative in front that cannot be ignored

hero (hero):

Are you sure you wrote down the problem correctly?

OpenStudy (anonymous):

k wait

OpenStudy (anonymous):

k it was sec not -sec sorry i just wrote it :- instead of writing it :

OpenStudy (anonymous):

can u plz help me with other one?

OpenStudy (anonymous):

cotx-cos^3xcscx

hero (hero):

\[\cot x - \cos^3x \csc x \\=\frac{\cos x}{\sin x} - \frac{\cos^3x}{\sin x} \\=\frac{\cos x - \cos^3x}{\sin x} \\=\frac{\cos x(1 - \cos^2x)}{\sin x} \\=\frac{\cos x(\sin^2x)}{\sin x} \\=\cos x \sin x \]

OpenStudy (anonymous):

thank you so much u rock!i wish u were my pre cal teacher

hero (hero):

okay

OpenStudy (anonymous):

wait how did u get cos^3x-cosx=cosx(1-cos^2x) ?

OpenStudy (anonymous):

can't we just subtract it out?

OpenStudy (anonymous):

k i get what u did !thanks for all ur help